RC meeting and Run Plan, Monday January 4th
From SBS wiki
Contents
RC Meeting Information
Run Coordinator: Bogdan Wojtsekhowski
Location: Counting House Conference Room
Time: 3:15 pm
Remote connection link: Zoomgov
Toll-Free Number (U.S.& Canada): 1-888-240-2560
Bluejeans Meeting ID: 256 205 934
Upcoming Shifts Which Need Filled
VERY URGENT: Shifts in January need help - 100+ openings to fill
Today's estimate: Optic beam Tue. Jan 11, "Bonus beam" Jan 12-16. Official beam Jan 17
Accelerator and Hall Status
- Preparation for beam start
To be Done During Extended Down Time
- Replace J0 INFN plane with UVa plane UV#4 and related electronics
- Spectrometers move to positions for SBS-14 kinematics
- Cosmic tests of HCal and BBcal
- Study of GRINCH
- Checks quality of the fiber connections within DAQ
Run Plan
=
- None
Resolved Issues
Unresolved Issues
- NEW: HCAL ROC17 ERROR for last 12+ hours
- Occasional timing shifts of ~70 ns
- GEM efficiency analysis.
Subsystem Updates
- DAQ general
- Alex update. Statistics of DAQ reboots.
- HCal
- F1 errors (HCal ROC17 sync errors)... Don't ignore!! End, Reset, Configure, Download. If that doesn't work, call DAQ on call.
- Could 50k replay show a problem FLAG?
- News on time resolution and amplitude analysis.
- BBCal
- Lesson from BBcal HV loss for few hours on Friday. Update on monitoring/reboot methods.
- Timing Hodoscope
- Update on TH setup and time resolution. Plan for spare parts.
- GRINCH
- Update from data analysis.
- GEMS
- DAQ update
- HV and efficiency status. Perspectives for higher luminosity.
- Beamline
- Update on raster analysis.
- Replay analysis
- Massive data replay: script preparation is in progress.
- Observations on data quality.